### v1

# On the birational invariance of the balanced hyperbolic manifolds

Abstract
In this paper, we discuss the birational invariance of the class of balanced hyperbolic manifolds.
32Q15, 32J25, 32J27
keywords:
Balanced metric, Hyperbolicity, Birational invariant
1 Introduction
In his celebrated paper [Gro91], M. Gromov introduces an important notion called the Kähler hyperbolicity. It is pinched between the real hyperbolicity and the Kobayashi hyperbolicity [Kob98], and helps to settle the Kähler case of the Chern conjecture [Gro91]. After that, it leads to fruitful applications and improvements such as [CX01, CY18, Kol95, Eys97, Hit00, McM00] and so on.
However, since the class of Kähler manifolds in general is not invariant under the birational transform, it would be desirable to have a birational variant of the Kähler hyperbolicity developed. It is an open problem posed by J. Kollár in [Kol95]. Kollár suggests to require Gromov’s condition for a degenerated Kähler form, and [BDET22, BCDT24] introduce the weakly Kähler hyperbolicity by asking the cohomology class to be nef and big rather than Kähler. Weakly Kähler hyperbolic manifolds possess many key features as Kähler hyperbolic manifolds, and are invariant under the birational transform. Whereas in this paper, we are trying to investigate a more general situation, namely the balanced hyperbolicity.
More precisely, let $X$ be a compact complex manifold of dimension $n$. A Hermitian metric $\omega$ on $X$ is called balanced if $d\omega^{n-1}=0$. $X$ is called a balanced manifold if it possesses a balanced metric. Obviously a Kähler metric must be balanced, but there do exist non-Kähler balanced metrics. Hence a balanced form is also regarded as a degenerated Kähler form in some aspects. A celebrated theorem in [AB95] asserts that the class of compact balanced manifolds is invariant under the birational transform, which directly inspires this paper.
Now, let $\pi:\tilde{X}\rightarrow X$ be the universal cover, and fix a Riemannian metric $g$ on $X$. Recall that a $k$-form $\alpha$ on $X$ is called $\tilde{d}$-bounded, if there exists a $(k-1)$-form $\beta$ on $\tilde{X}$ such that $\pi^{\ast}\alpha=d\beta$ and $\sup_{\tilde{X}}\|\beta\|_{\pi^{\ast}g}<\infty$. Note since $X$ is compact, this notion is actually independent of the choice of $g$. We say that a Hermitian metric $\omega$ on $X$ is balanced hyperbolic if $\omega$ is balanced and $\omega^{n-1}$ is $\tilde{d}$-bounded. It is notable that the balanced hyperbolicity was first introduced in [MP22, MP23], and we will continue this streamline to make a discussion in a wider range.
Observe that if $\omega$ is balanced, $[\omega^{n-1}]$ must be nef and big as an $(n-1,n-1)$-class. It allows us to talk about more degenerated cases. We say that a real smooth $(1,1)$-form $\omega$ is semi-balanced hyperbolic if $\omega^{n-1}$ is $d$-closed, non-negative, strictly positive on a Zariski open set, and $\tilde{d}$-bounded. Fix an integer $k\mid(n-1)$, say $kt=n-1$. We say that a real smooth $(k,k)$-form $\beta$ is weakly balanced $k$-hyperbolic if $\beta^{t}$ is $d$-closed, $\beta$ is $\tilde{d}$-bounded and $[\beta^{t}]$ is nef and big. Finally, we say $X$ is a semi-balanced hyperbolic (resp. weakly balanced $k$-hyperbolic) manifold, if there exists a semi-balanced hyperbolic (resp. weakly balanced $k$-hyperbolic) form on it.
In a similar atmosphere, we can also talk about the $k$-hyperbolicity for the Kähler case. We say that a Hermitian metric $\omega$ on $X$ is Kähler $k$-hyperbolic if $\omega$ is Kähler and $\omega^{k}$ is $\tilde{d}$-bounded. We say that a real smooth $(1,1)$-form $\omega$ is semi-Kähler $k$-hyperbolic if $\omega$ is $d$-closed, non-negative, strictly positive on a Zariski open set and $\omega^{k}$ is $\tilde{d}$-bounded. We say that a real smooth $(1,1)$-form $\omega$ is weakly Kähler $k$-hyperbolic if $\omega$ is $d$-closed, $[\omega]$ is nef and big, and $\omega^{k}$ is $\tilde{d}$-bounded. Finally, we say $X$ is a (weakly or semi-) Kähler $k$-hyperbolic manifold, if there exists a (weakly or semi-) Kähler $k$-hyperbolic form on it. In Sect. 2.2 we will discuss the relationship among these hyperbolicities.
Then we can state our first result concerning the behavior of the balanced hyperbolicity under the small contraction. Note throughout this paper, a small contraction $f:\hat{X}\rightarrow X$ between compact complex manifolds is a holomorphic map such that, for a suitable analytic variety $Y$ with $\textrm{codim}Y\geqslant 2$ in $X$, $E:=f^{-1}(Y)$ (called the exceptional set) is codimensional $\geqslant 2$ and $f|_{\hat{X}\setminus E}:\hat{X}\setminus E\rightarrow X\setminus Y$ is a biholomorphism.
Theorem 1.1.
Let $f:\hat{X}\rightarrow X$ be a small contraction between compact complex manifolds of dimension $n$.
  1. (1)
    Assume that the higher homotopy groups $\pi_{i}(\hat{X})=0$ for $32\leqslant i\leqslant 2n-$ 3. Then $\hat{X}$ is balanced hyperbolic if and only if $X$ is.
  2. (2)
    If $X$ is semi-balanced hyperbolic, so will be $\hat{X}$.
  3. (3)
    Fix a positive integer $k\mid(n-1)$. Assume that both $\hat{X}$ and $X$ are Kähler, and the higher homotopy groups $\pi_{i}(\hat{X})=0$ for $12\leqslant i\leqslant 2k-$ 1. If $\hat{X}$ is weakly balanced $k$-hyperbolic, so will be $X$.
The only purpose for the Kähler assumption in (3) is to establish the cone duality between certain $(1,1)$ and $(n-1,n-1)$-classes as we will see later.
Next, we investigate the smooth modification. A smooth modification between compact complex manifolds $f:\hat{X}\rightarrow X$ in our paper is almost the same as the small contraction, except that $E$ is now a hypersurface. In this setting, we obtain the following results.
Theorem 1.2.
Let $f:\hat{X}\rightarrow X$ be a smooth modification between compact complex manifolds of dimension $n$.
  1. (1)
    Assume that the higher homotopy groups $\pi_{i}(\hat{X})=0$ for $32\leqslant i\leqslant 2n-$ 3. If $\hat{X}$ is balanced hyperbolic, so will be $X$.
  2. (2)
    If $X$ is semi-balanced hyperbolic, so will be $\hat{X}$.
  3. (3)
    Fix a positive integer $k\mid(n-1)$. Assume that both $\hat{X}$ and $X$ are Kähler, and the higher homotopy groups $\pi_{i}(\hat{X})=0$ for $12\leqslant i\leqslant 2k-$ 1. If $\hat{X}$ is weakly balanced $k$-hyperbolic, so will be $X$.
Remark 1.1.
[BCDT24] develops a beautiful topological technique to show that $\hat{X}$ is weakly Kähler 1-hyperbolic if and only if $X$ is. Our Theorems 1.1 and 1.2 owe it a lot.
In the ending section, we attempt to get rid of the vanishing condition about the higher homotopy groups. It is possible at least when $\dim E=1$, i.e. $f$ is a curve contraction.
Theorem 1.3.
Let $f:\hat{X}\rightarrow X$ be a curve contraction between compact complex manifolds of dimension $n$.
  1. (1)
    If $\dim X=2$ and $\hat{X}$ is balanced hyperbolic, so will be $X$.
  2. (2)
    If $\dim X>2$, then $\hat{X}$ is balanced hyperbolic if and only if $X$ is.
  3. (3)
    If $\hat{X}$ is weakly balanced $(n-1)$-hyperbolic, so will be $X$.
It naturally leads to the following generalization.
Theorem 1.4.
Fix a positive integer $k\mid(n-1)$. Let $f:\hat{X}\rightarrow X$ be a birational transform between compact Kähler manifolds of dimension $n$, such that the exceptional set $E$ satisfies $\dim E=n-k$. If $\hat{X}$ is a weakly balanced $k$-hyperbolic, so will be $X$.
We call both small contraction and smooth modification the birational transform in this paper. All of the theorems above show that these hyperbolicities surely possess certain invariance under the birational transform, and are also good candidates to solve Kollár’s problem.
2 Preliminary
Let $X$ be a compact complex manifold of dimension $n$.
2.1 Positive cones
We first recall a few positive cones in the Bott–Chern cohomology groups
$H^{1,1}_{\textrm{BC}}(X,\mathbb{R})\quad\textrm{and}\quad H^{n-1,n-1}_{\textrm{BC}}(X,\mathbb{R}).$
Note a $(1,1)$-class is usually called the divisor class, while an $(n-1,n-1)$-class the curve class.
Definition 2.1.
  1. 1.
    Let
    $\mathcal{E}(X):=\{[T]\in H^{1,1}_{\textrm{BC}}(X,\mathbb{R})|T\textrm{ is a positive }d\textrm{-closed }(1,1)\textrm{-current on }X\},$
    and
    $\mathcal{N}(X):=\{[T]\in H^{n-1,n-1}_{\textrm{BC}}(X,\mathbb{R})|T\textrm{ is a positive }d\textrm{-closed }(n-1,n-1)\textrm{-current on }X\}.$
    They are obviously closed cones in $H^{1,1}_{\textrm{BC}}(X,\mathbb{R})$ and $H^{n-1,n-1}_{\textrm{BC}}(X,\mathbb{R})$ respectively, which are called the pseudo-effective divisor cone and the pseudo-effective curve cone.
  2. 2.
    $\mathcal{E}(X)^{\circ}$ (resp. $\mathcal{N}(X)^{\circ}$) is called the big divisor (resp. curve) cone.
  3. 3.
    Let
    $\mathcal{K}(X):=\{[\omega]\in H^{1,1}_{\textrm{BC}}(X,\mathbb{R})|\omega\textrm{ is a smooth, strictly positive }d\textrm{-closed }(1,1)\textrm{-form on }X\},$
    and
    $\begin{split}\mathcal{B}(X):=&\{[\Omega]\in H^{n-1,n-1}_{\textrm{BC}}(X,\mathbb{R})|\Omega\textrm{ is a smooth, strictly positive }d\textrm{-closed }\\ &(n-1,n-1)\textrm{-form on }X\}.\end{split}$
    They are obviously open cones in $H^{1,1}_{\textrm{BC}}(X,\mathbb{R})$ and $H^{n-1,n-1}_{\textrm{BC}}(X,\mathbb{R})$ respectively, which are called the Kähler cone and the balanced cone.
  4. 4.
    Fix a Hermitian metric $\sigma$ on $X$. Let
    $\begin{split}\overline{\mathcal{K}(X)}:=&\{[\omega]\in H^{1,1}_{\textrm{BC}}(X,\mathbb{R})|\textrm{For any }\varepsilon>0,\textrm{there exists a smooth representative }\omega_{\varepsilon}\in[\omega]\\ &\textrm{ such that }\omega_{\varepsilon}\geqslant-\varepsilon\sigma\},\end{split}$
    and
    $\begin{split}\overline{\mathcal{B}(X)}:=&\{[\Omega]\in H^{n-1,n-1}_{\textrm{BC}}(X,\mathbb{R})|\textrm{For any }\varepsilon>0,\textrm{there exists a smooth representative }\\ &\Omega_{\varepsilon}\in[\Omega]\textrm{ such that }\Omega_{\varepsilon}\geqslant-\varepsilon\sigma^{n-1}\}.\end{split}$
    $\overline{\mathcal{K}(X)}$ (resp. $\overline{\mathcal{B}(X)}$) is called the nef divisor (resp. curve) cone. Clearly, $\overline{\mathcal{K}(X)}\subseteq\mathcal{E}(X)$ and $\overline{\mathcal{B}(X)}\subseteq\mathcal{N}(X)$. When $\mathcal{K}(X)$ (resp. $\mathcal{B}(X)$) is not empty, $\overline{\mathcal{K}(X)}$ (resp. $\overline{\mathcal{B}(X)}$) is its closure.
A $(1,1)$-class is called pseudo-effective (resp. big, nef,…) if it belongs to $\mathcal{E}(X)$ (resp. $\mathcal{E}(X)^{\circ}$, $\overline{\mathcal{K}(X)}$,…). The positivity for the $(n-1,n-1)$-classes is reformulated similarly.
We should recall some basic properties of the balanced cone in [FX14]. Remember that a Hermitian metric $\omega$ such that $d\omega^{n-1}=0$ is called a balanced metric. We say $X$ is a balanced manifold, if there exists a balanced metric on it. It is easy to see that the existence of a balanced metric $\omega$ is equivalent to the existence of a $d$-closed strictly positive $(n-1,n-1)$-form $\Omega$ with the relations
$\Omega=\omega^{n-1}\quad\textrm{and}\quad\omega=\frac{\ast\Omega}{(n-1)!}.$
Hence, for convenience, each such $\Omega$ will also be called a balanced metric, which leads to the definition of the balanced cone $\mathcal{B}(X)$ above. In this paper, we will always use $\omega,\hat{\omega}$ to denote the $(1,1)$-form and the capital letters such as $\Omega,\hat{\Omega}$ to denote the $(n-1,n-1)$-form.
The relationship between the balanced cone and the divisor cone is subtle. We denote by $A^{p,q}(X)$ the space of the smooth $\mathbb{C}$-valued $(p,q)$-forms and by $A^{p,q}_{\mathbb{R}}(X)$ the space of the smooth $\mathbb{R}$-valued $(p,q)$-forms, then
$V^{p,p}(X,\mathbb{R}):=\frac{\{\phi\in A^{p,p}_{\mathbb{R}}(X)|\partial\bar{\partial}\phi=0\}}{\{\partial A^{p-1,p}(X)+\bar{\partial}A^{p,p-1}(X)\}\cap A^{p,p}_{\mathbb{R}}(X)}.$
It is well-known that we can replace $A^{p,p}$ by the space of $(p,p)$-currents in the above definition. If we denote ¯$\mathcal{E}_{\partial\bar{\partial}}(X)\subseteq V^{1,1}(X,\mathbb{R})$ the convex cone generated by $\partial\bar{\partial}$-closed positive $(1,1)$-currents, it is proved in [FX14] (Lemma 3.3 & Remark 3.4) that
Lemma 2.1.
¯$\mathcal{E}_{\partial\bar{\partial}}(X)^{\vee}=\overline{\mathcal{B}(X)}$.
Then combining with a celebrated description for a balanced manifold, we reformulate it as
Proposition 2.1.
Let $X$ be a compact balanced manifold of dimension $n$, and let $\Omega$ be a real $d$-closed $(n-1,n-1)$-form on $X$. Then the following two statements are equivalent:
  1. (a)
    $[\Omega]$ is a balanced class;
  2. (b)
    for any positive $\partial\bar{\partial}$-closed $(1,1)$-current $T$, $\int_{X}\Omega\wedge T\geqslant 0$ and $\int_{X}\Omega\wedge T=0$ if and only if $T=0$.
Proof.
(b)$\Rightarrow$(a) is nothing but [FX14], Lemma 3.3.
(a)$\Rightarrow$(b) is due to [Mic83], Proposition 4.2. ∎
Remark 2.1.
When $V^{1,1}(X,\mathbb{R})=H^{1,1}_{\textrm{BC}}(X,\mathbb{R})$, say $X$ is a Kähler manifold, Lemma 2.1 is reformulated as
$\mathcal{E}(X)^{\vee}=\overline{\mathcal{B}(X)}.$ (2.1)
2.2 The hyperbolicity
With the preparation above, we are able to define various hyperbolicities mentioned in Introduction.
Definition 2.2.
For $k=1,...,n$, we say that a Hermitian metric $\omega$ is balanced $k$-hyperbolic, if it is balanced and $\omega^{k}$ is $\tilde{d}$-bounded. We say that a real smooth $(1,1)$-form $\omega$ is semi-balanced $k$-hyperbolic, if $\omega^{n-1}$ is $d$-closed, non-negative, strictly positive on a Zariski open set and $\omega^{k}$ is $\tilde{d}$-bounded. Fix an integer $k\mid(n-1)$, say $kt=n-1$. We say that a real smooth $(k,k)$-form $\beta$ is weakly balanced $k$-hyperbolic, if $\beta^{t}$ is $d$-closed, $[\beta^{t}]$ is nef and big, and $\beta$ is $\tilde{d}$-bounded.
Let $X$ be a balanced manifold. We say $X$ is a (weakly or semi-) balanced $k$-hyperbolic manifold, if there exists a (weakly or semi-) balanced $k$-hyperbolic form on it.
Remember that a Hermitian metric $\omega$ is Kähler hyperbolic [Gro91] if it is Kähler and $\tilde{d}$-bounded; whereas a real smooth $(1,1)$-form $\omega$ is semi-Kähler hyperbolic [Kol95], if $\omega$ is $d$-closed, non-negative, strictly positive on a Zariski open set and $\tilde{d}$-bounded; whereas a real smooth $(1,1)$-form $\omega$ is weakly Kähler hyperbolic [BDET22], if it is $d$-closed, $\tilde{d}$-bounded, and $[\omega]$ is nef and big.
We can also generalize these notions to the $k$-hyperbolicity.
Definition 2.3.
For $k=1,...,n$, we say that a Hermitian metric $\omega$ is Kähler $k$-hyperbolic, if it is Kähler and $\omega^{k}$ is $\tilde{d}$-bounded. We say that a real smooth $(1,1)$-form $\omega$ is semi-Kähler $k$-hyperbolic, if $\omega$ is $d$-closed, non-negative, strictly positive on a Zariski open set and $\omega^{k}$ is $\tilde{d}$-bounded. We say that a real smooth $(1,1)$-form $\omega$ is weakly Kähler $k$-hyperbolic, if $\omega$ is $d$-closed, $[\omega]$ is nef and big, and $\omega^{k}$ is $\tilde{d}$-bounded.
Let $X$ be a Kähler manifold. We say $X$ is a (weakly or semi-) Kähler $k$-hyperbolic manifold, if there exists a (weakly or semi-) Kähler $k$-hyperbolic form on it.
Next we make a discussion about the relationship among these hyperbolicities. Clearly, Gromov’s (weakly or semi-)Kähler hyperbolicity is nothing but (weakly or semi-)Kähler $1$-hyperbolicity in our language. For a balanced $k$-hyperbolic metric $\omega$, $\pi^{\ast}d\omega^{k}=d\pi^{\ast}\omega^{k}=0$ since $\pi^{\ast}\omega^{k}$ is $d$-exact. However, it implies that $d\omega^{k}=0$ because $\tilde{X}\rightarrow X$ is locally isomorphic. Moreover, when $k<n-1$, $d\omega^{k}=0$ indicates that $d\omega=0$ by standard multi-linear algebra. All in all, we conclude that a balanced $k$-hyperbolic manifold with $k<n-1$ must be Kähler $k$-hyperbolic. It also happens when $\omega$ is semi-balanced. We first obtain $d\omega=0$ on the Zariski open set where $\omega$ is strictly positive, then deduce that $d\omega=0$ on the whole space by smoothness. Hence the (semi-)balanced $k$-hyperbolicity only makes sense when $k=n-1$, in which case we simply call it (semi-)balanced hyperbolicity. Clearly this argument fails when $\omega$ further degenerates.
By definition, a Kähler $k$-hyperbolic manifold must be both semi-Kähler $k$-hyperbolic and balanced $k$-hyperbolic with $k=1,...,n$. A semi-Kähler $k$-hyperbolic manifold is clearly both weakly Kähler $k$-hyperbolic and semi-balanced $k$-hyperbolic. A balanced hyperbolic manifold must be semi-balanced hyperbolic. Also we have
Lemma 2.2.
A weakly Kähler $k$-hyperbolic manifold must be weakly balanced $k$-hyperbolic for $k=1,...,n$.
Proof.
Let’s recall two crucial facts in [LX16]. Let $[\alpha]\in\mathcal{N}(X)$, and define the volume to be
$\widehat{\textrm{vol}}([\alpha])=\inf_{[A]\textrm{ big and nef (1,1)-class}}(\frac{[A]\cdot[\alpha]}{\textrm{vol}([A])^{1/n}})^{\frac{n}{n-1}}.$
Then [LX16] indicates that
Theorem 2.1 ((c.f. [LX16], Theorem 5.2)).
  1. (1)
    $\widehat{\textrm{vol}}$ is positive precisely for the big classes.
  2. (2)
    For any big and nef $(1,1)$-class $[A]$, we have $\widehat{\textrm{vol}}([A^{n-1}])=\textrm{vol}([A])$.
Although the whole article is reformulated for projective manifolds, it is also carefully explained in [LX16], Sect.2.4 that everything extends smoothly to the Kähler case.
Now if $X$ is a weakly Kähler $k$-hyperbolic manifold, then there exists a nef and big $(1,1)$-class $[\omega]$ such that $\omega^{k}$ is $\tilde{d}$-bounded. So $\widehat{\textrm{vol}}([\omega^{n-1}])=\textrm{vol}([\omega])>0$. The inequality is due to [DP04], Theorem 0.5. Therefore $[\omega^{n-1}]$ is a nef and big $(n-1,n-1)$-class. It exactly implies that $X$ is weakly balanced $k$-hyperbolic. ∎
For a balanced metric $\omega$, $\omega^{n-1}$ is certainly positive hence $[\omega^{n-1}]$ is nef. It means that $\mathcal{B}(X)$ is an open subcone of $\mathcal{N}(X)$, so $\mathcal{B}(X)\subseteq\mathcal{N}(X)^{\circ}$. Therefore $[\omega^{n-1}]$ is also big. It means that a balanced hyperbolic metric must be weakly balanced hyperbolic.
In the end, let $\pi:\tilde{X}\rightarrow X$ be the universal cover, and fix a Riemannnian metric $g$ on $X$. If $\omega^{k}$ is $\tilde{d}$-bounded, then there exists a $(2k-1)$-form $\eta$ on $\tilde{X}$ such that $\pi^{\ast}\omega^{k}=d\eta$ and $\eta$ is bounded with respect to $\pi^{\ast}g$. Then for any integer $t$,
$\pi^{\ast}\omega^{kt}=d\eta\wedge\pi^{\ast}\omega^{k(t-1)}=d(\eta\wedge\pi^{\ast}\omega^{k(t-1)}).$
Since $X$ is compact, $\omega$ is bounded with respect to $g$. Hence $\pi^{\ast}\omega^{k(t-1)}$ as well as $\eta\wedge\pi^{\ast}\omega^{k(t-1)}$ is bounded with respect to $\pi^{\ast}g$. It exactly means that the $k$-hyperbolicity implies the $kt$-hyperbolicity for any positive integer $k,t$.
We summarize the relationship among these notions as follows: for any positive integers $k,t$,
[…]
and
[…]
It worthwhile to point out that there exist weakly Kähler but not Kähler 1-hyperbolic manifolds provided by [BDET22, BCDT24]. Moreover, the connected sums
#$\#_{k}(S^{3}\times S^{3})\textrm{ with }k\geqslant 2$
are balanced manifolds as is shown in [FLY12]. A direct computation implies that
$H^{4}_{\textrm{dR}}(\#_{k}(S^{3}\times S^{3}),\mathbb{R})=0.$
So they cannot be compact Kähler manifolds. On the other hand, any balance metric $\omega$ gives a zero class $[\omega^{2}]$. It means that $\omega$ is balanced hyperbolic. Therefore there exist non-Kähler (weakly or semi-)balanced manifolds, and the horizontal implications in the diagram are strict except the mentioned cases. We are willing to know more information about the vertical inclusions. For example, must a semi-balanced hyperbolic manifold be weakly balanced $(n-1)$-hyperbolic?
We also have the following interesting description for the fundamental group of a weakly balanced hyperbolic manifold, though it is not really involved in our main theorems.
Proposition 2.2.
Let $X$ be a compact Kähler manifold. If $k\mid(n-1)$, and $X$ is weakly balanced $k$-hyperbolic, then $\pi_{1}(X)$ is not amenable.
Proof.
Let $\pi:\tilde{X}\rightarrow X$ be the universal cover. Denote by $H^{i}_{\beta}(\tilde{X})$ the de Rham cohomology based on differential forms $\alpha$ such that $\alpha$ and $d\alpha$ are uniformly bounded. Now assume that $\pi_{1}(X)$ is amenable. Due to [ABW92], the pull-back homomorphism $H^{i}(X)\rightarrow H^{i}_{\beta}(\tilde{X})$ is injective. But this is impossible.
In fact, the pull-back of the weakly balanced $k$-hyperbolic class $[\gamma]$ is zero in $H^{2k}_{\beta}(\tilde{X})$. By hypothesis $k\mid(n-1)$, say $kt=n-1$. So $\pi^{\ast}[\gamma^{t}]$ is also a zero class in $H^{2n-2}_{\beta}(\tilde{X})$. However, $[\gamma^{t}]$ is big hence never a zero class in $H^{2n-2}(X)$. Otherwise, $[A]\cdot[\gamma^{t}]=0$ for any $(1,1)$-class $[A]$. Hence $\widehat{\textrm{vol}}([\gamma^{t}])=0$, which leads to a contradiction by Theorem 2.1. All in all, $\pi_{1}(X)$ is not amenable. ∎
To the end of this section, we provide the following property that is frequently used when verifying the $\tilde{d}$-boundedness.
Lemma 2.3.
Let $f:\hat{X}\rightarrow X$ be a birational transform between compact complex manifolds. Fix Riemannian metrics $\hat{g}_{X},g_{X}$ on $\hat{X},X$ respectively, and suppose that $\omega$ is a $\tilde{d}$-bounded $k$-form on $X$. Let $\xi$ be an arbitrary smooth $(k-1)$-form on $X$. Then both $f^{\ast}\omega$ and $\omega+d\xi$ are $\tilde{d}$-bounded.
Proof.
Let $\pi:\tilde{X}\rightarrow X$ be the universal cover. Since $f$ is birational,
$f_{\ast}:\pi_{1}(\hat{X})\rightarrow\pi_{1}(X)$
is an isomorphism (see [BP21], Proposition 2.3). Therefore $\chi:Z:=f^{\ast}\tilde{X}\rightarrow\hat{X}$ gives a universal cover of $\hat{X}$. Denote by $h$ the natural morphism $Z\rightarrow\tilde{X}$.
Since $\omega$ is $\tilde{d}$-bounded, there exists a bounded $(k-1)$-form $\eta$ on $\tilde{X}$ such that $\pi^{\ast}\omega=d\eta$. Now by [BDET22], Lemma 2.28, $h^{\ast}\eta$ is also bounded with respect to $h^{\ast}\hat{g}_{X}$. Obviously $\chi^{\ast}f^{\ast}\omega=dh^{\ast}\eta$, so $f^{\ast}\omega$ is $\tilde{d}$-bounded.
The $\tilde{d}$-boundedness of $\omega+d\xi$ is much easier. Since $X$ is compact, $\xi$ is bounded with respect to $g_{X}$. Then $\eta+\pi^{\ast}\xi$ is also bounded with respect to $\pi^{\ast}g_{X}$, and $\pi^{\ast}(\omega+d\xi)=d(\eta+\pi^{\ast}\xi)$. So $\omega+d\xi$ is $\tilde{d}$-bounded. ∎
3 The birational invariance
This section is devoted to prove Theorems 1.1 and 1.2. We will see that the ingredient is a topological technique developed in [BCDT24].
3.1 Topological preparation I
Let $X$ be a simplicial complex. Then a $k$-form $\omega$ on $X$ consists of a smooth $k$-form $\omega_{\sigma}$ for every simplex $\sigma\subseteq X$ such that $\omega_{\sigma}|_{\tau}\equiv\omega_{\tau}$ whenever $\tau\subseteq\sigma$ is a subsimplex. A Riemannian metric on $X$ is a choice of a Riemannian metric $g_{\sigma}$ on every simplex $\sigma\subseteq X$ such that $g_{\sigma}|_{\tau}\equiv g_{\tau}$ for $\tau\subseteq\sigma$. Under these conventions we are able to talk about the $\tilde{d}$-bounded class on $X$.
We should recall and generalize the descriptions on hyperbolic classes in [BKS24, BCDT24]. Let $V^{k}_{\textrm{asph}}(X)$ be the subspace of $H^{k}(X,\mathbb{R})$, which consists of the $k$-cohomology classes whose pull-back to any real $k$-dimensional sphere is zero. Let $V^{k}_{\textrm{hyp}}(X)$ be the subspace of $H^{k}(X,\mathbb{R})$, which consists of the $\tilde{d}$-bounded $k$-cohomology classes. We will say an element $[\omega]\in V^{k}_{\textrm{asph}}(X)$ (resp. $[\omega]\in V^{k}_{\textrm{hyp}}(X)$) is aspherical (resp. hyperbolic). As soon as $k\geqslant 2$, every continuous map from a $k$-dimensional sphere $S^{k}\rightarrow X$ factorizes through the universal cover of $X$, so $V^{k}_{\textrm{hyp}}(X)\subseteq V^{k}_{\textrm{asph}}(X)$ for $k\geqslant 2$.
Let $G=\pi_{1}(X)$ and consider the classifying space $EG\rightarrow BG$. Given a universal cover $\tilde{X}\rightarrow X$, there is a unique (up to homotopy) classifying map of this universal cover
$c_{\tilde{X},X}:X\rightarrow BG$
such that $\tilde{X}$ is isomorphic to the pull-back $c^{\ast}_{\tilde{X},X}EG$ as a $G$-principal bundle. Define the subspace $V^{k}_{\textrm{hyp}}(BG)$ of the real singular cohomology group $H^{k}(BG,\mathbb{R})$ to be the set of $k$-cohomology classes whose pull-back to any finite simplicial complex is hyperbolic.
Then we have
Theorem 3.1 ((c.f. [BKS24], Theorem 2.5)).
Fix an integer $k\geqslant 2$. Let $X,Y$ be two finite simplicial complexes such that $\pi_{i}(X)=0$ for $12\leqslant i\leqslant k-$ 1. Let $f:Y\rightarrow BG$ be an arbitrary continuous map, and let $[\omega]\in H^{k}(BG,\mathbb{R})$. If $c^{\ast}_{\tilde{X},X}[\omega]\in V^{k}_{\textrm{hyp}}(X)$, then $f^{\ast}[\omega]\in V^{k}_{\textrm{hyp}}(Y)$.
In particular, Theorem 3.1 implies that if $c^{\ast}_{\tilde{X},X}[\omega]$ is hyperbolic, so is $[\omega]$. Also we have
Lemma 3.1.
Fix an integer $k\geqslant 2$. Then
$c^{\ast}_{\tilde{X},X}(H^{k}(BG,\mathbb{R}))=V^{k}_{\textrm{asph}}(X).$
If we furthermore assume that $\pi_{i}(X)=0$ for $2\leqslant i\leqslant k$, the pull-back
$c^{\ast}_{\tilde{X},X}:H^{k}(BG,\mathbb{R})\rightarrow H^{k}(X,\mathbb{R})$
is even injective.
Proof.
Firstly, every class in $H^{k}(BG,\mathbb{R})$ is aspherical as is pointed out in [BCDT24], Lemma 2.9. On the other hand, the pull-back of an aspherical class via $c^{\ast}_{\tilde{X},X}$ is still aspherical by definition. It implies
$c^{\ast}_{\tilde{X},X}(H^{k}(BG,\mathbb{R}))=c^{\ast}_{\tilde{X},X}(V^{k}_{\textrm{asph}}(BG))\subseteq V^{k}_{\textrm{asph}}(X).$
Next we construct a model of $BG$ by attaching cells of real dimension $3$ or higher to $X$ to make the universal cover contractible without affecting $\pi_{1}(X)$. In this situation, $X\subseteq BG$, and the classifying map $c_{\tilde{X},X}$ is the inclusion. We now from the long exact sequence of the pair $X\subseteq BG$ in cohomology, to get
[…]
It remains to prove that if $[\omega]\in V^{k}_{\textrm{asph}}(X)$, then $\delta[\omega]=0$. In fact, by definition
$\delta[\omega]\in H_{k+1}(BG,X,\mathbb{R})^{\ast}$
acts on a relative $(k+1)$-cycle $\gamma+C_{k+1}(X)$ by $\delta[\omega](\gamma+C_{k+1}(X))=\omega(\partial_{k+1}\gamma)$, where $\partial_{k+1}$ is the boundary operator. Observe that $\partial_{k+1}\gamma\in Z_{k}(X)$ is a $k$-cycle in $X$ which, if non-trivial, comes from a $(k+1)$-cell attached to $X$. It means that $\partial_{k+1}\gamma$ is linear combinations of images of $S^{k}$. Therefore, $\omega(\partial_{k+1}\gamma)=0$ since $[\omega]$ is aspherical.
In the end, if $\pi_{i}(X)=0$ for $2\leqslant i\leqslant k$, the relative homology group $H_{k}(BG,X,\mathbb{R})$ vanishes by the relative Hurewicz theorem. So does $H^{k}(BG,X,\mathbb{R})$ by duality. Then we obtain the injectivity from the exact sequence above immediately. ∎
Combining with Theorem 3.1 and Lemma 3.1, we obtain that
Corollary 3.1.
Fix an integer $k\geqslant 2$. Let $X$ be a finite simplicial complex such that $\pi_{i}(X)=0$ for $12\leqslant i\leqslant k-$ 1. Then
$c^{\ast}_{\tilde{X},X}(V^{k}_{\textrm{hyp}}(BG))=V^{k}_{\textrm{hyp}}(X).$
Proof.
By definition $c^{\ast}_{\tilde{X},X}(V^{k}_{\textrm{hyp}}(BG))\subseteq V^{k}_{\textrm{hyp}}(X)$. For the opposite inclusion, if $[\alpha]\in V^{k}_{\textrm{hyp}}(X)$, then $[\alpha]\in V^{k}_{\textrm{asph}}(X)$, and thus by Lemma 3.1 there exists $[\omega]\in H^{k}(BG,\mathbb{R})$ such that $[\alpha]=c^{\ast}_{\tilde{X},X}[\omega]$. So $c^{\ast}_{\tilde{X},X}[\omega]$ is hyperbolic. But then, by Theorem 3.1, $[\omega]\in V^{k}_{\textrm{hyp}}(BG)$. ∎
3.2 The small contraction
As a warm-up, we first prove the class of balanced hyperbolic manifolds is invariant under the small contraction $f:\hat{X}\rightarrow X$ between compact complex manifolds of dimension $n$. The ingredient is the birational invariance of the balanced metric proved in [AB91, AB92, AB95], which we sketch as follows.
One direction is simple. Assume that $(X,\omega)$ is balanced. By Proposition 2.1, it means that for any positive $\partial\bar{\partial}$-closed $(1,1)$-current $T$ on $X$, $\int_{X}\omega^{n-1}\wedge T\geqslant 0$ and $\int_{X}\omega^{n-1}\wedge T=0$ if and only if $T=0$. Then we consider $\hat{\omega}:=f^{\ast}\omega$. For any positive $\partial\bar{\partial}$-closed $(1,1)$-current $S$ on $\hat{X}$, a direct computation implies that
$\int_{\hat{X}}f^{\ast}\omega^{n-1}\wedge S=\int_{X}\omega^{n-1}\wedge f_{\ast}S\geqslant 0.$
Moreover, $\int_{\hat{X}}f^{\ast}\omega^{n-1}\wedge S=0$ if and only if $f_{\ast}S=0$. It means that $\textrm{supp}S\subseteq E$. Since $\textrm{codim}E\geqslant 2$, we can apply the following proposition to deduce that $S=0$. It implies that $[f^{\ast}\omega^{n-1}]$ is a balanced class on $\hat{X}$ by Proposition 2.1.
Proposition 3.1 ((c.f. [AB92], Theorem 1.1)).
Let $X$ be a complex manifold of dimension $n$. Assume $T$ is a $\partial\bar{\partial}$-closed positive $(p,p)$-current on $X$ such that the Hausdorff $2(n-p)$-measure of $\textrm{supp}T$ vanishes. Then $T=0$.
The converse is more involved. We will mainly use the following result, which is a simple variant of [AB95].
Theorem 3.2 ((c.f. [AB95], Theorem 3)).
Let $T$ be a positive $\partial\bar{\partial}$-closed $(1,1)$-current on $X$. Then there exists a unique positive $\partial\bar{\partial}$-closed $(1,1)$-current $\hat{T}$ on $\hat{X}$ such that $f_{\ast}\hat{T}=T$ and $\hat{T}\in f^{\ast}[T]\in V^{1,1}(\hat{X},\mathbb{R})$.
Proof.
Note the original Theorem 3 of [AB95] is reformulated for a smooth modification, i.e. $\textrm{codim}E=1$. However, everything also works well when $\textrm{codim}E>1$. In fact, the authors first prove this result when $f:\hat{X}\rightarrow X$ is a finite sequence of blow-ups with smooth center in [AB93]. Then with the help of the following Lemma 3.2, the desired result for a general smooth modification is standard.
Lemma 3.2 ((c.f. [AB93], Lemma 2.6)).
Let $f:\hat{X}\rightarrow X$ be a smooth modification. For every $x\in X$ there exists an open neighborhood $V$ of $x$ in $X$, a complex manifold $Z$ and holomorphic maps $g:Z\rightarrow\hat{X}$, $q:Z\rightarrow V$ such that $q=f\circ g$; moreover, $g:Z\rightarrow f^{-1}(V)$ is a blow-up and $q:Z\rightarrow V$ is obtained as a finite sequence of blow-ups with smooth center.
So it is left to check this lemma for a small contraction $f:\hat{X}\rightarrow X$. This is well-known to the experts. For readers’ benefit, one could applies verbatim the original proof of Lemma 3.2. Notice that the main tools there including Lemmas 4,7,8 and Corollary 1 of [HR64] are always applicable as long as $X\setminus Y$ (resp. $\hat{X}\setminus E$) is dense in $X$ (resp. $\hat{X}$), and $f|_{\hat{X}\setminus E}:\hat{X}\setminus E\rightarrow X\setminus Y$ is a biholomorphism. The concrete dimension of $E$ is not essential. ∎
Now assume that $(\hat{X},\hat{\Omega})$ is balanced. By Proposition 2.1, it means that for any positive $\partial\bar{\partial}$-closed $(1,1)$-current $S$ on $\hat{X}$, $\int_{\hat{X}}\hat{\Omega}\wedge S\geqslant 0$ and $\int_{\hat{X}}\hat{\Omega}\wedge S=0$ if and only if $S=0$. Then we consider $\Omega:=f_{\ast}\hat{\Omega}$, which is a $d$-closed $(n-1,n-1)$-current on $X$. For any positive $\partial\bar{\partial}$-closed $(1,1)$-current $T$ on $X$, a direct computation shows that
$\int_{X}\Omega\wedge T=\int_{\hat{X}}\hat{\Omega}\wedge\hat{T}\geqslant 0.$
Here $\hat{T}$ is the unique positive $\partial\bar{\partial}$-closed $(1,1)$-current on $\hat{X}$ obtained by Theorem 3.2. Moreover, $\int_{X}\Omega\wedge T=0$ if and only if $\hat{T}=0$. In this situation, $T=f_{\ast}\hat{T}=0$. It exactly implies that $[\Omega]$ is a balanced class on $X$.
Now we are ready to prove the birational invariance of the balanced hyperbolic manifolds.
Proof of Theorem 1.1.
(1) Assume that $(X,\omega)$ is balanced hyperbolic. As is shown above, the real $d$-closed $(n-1,n-1)$-class $f^{\ast}[\omega^{n-1}]$ is balanced. Hence there exists a balanced metric $\hat{\omega}$ such that $[\hat{\omega}^{n-1}]=f^{\ast}[\omega^{n-1}]$. On the other hand, since $\omega^{n-1}$ is $\tilde{d}$-bounded, so is $f^{\ast}\omega^{n-1}$ by Lemma 2.3. It implies that $\hat{\omega}^{n-1}$ is $\tilde{d}$-bounded again by Lemma 2.3.
Conversely, assume that $(\hat{X},\hat{\omega})$ is balanced hyperbolic. Let $G=\pi_{1}(X)$, and fix a classifying space $EG\rightarrow BG$. As is shown in Lemma 2.3, $\chi:Z:=f^{\ast}\tilde{X}\rightarrow\hat{X}$ gives a universal cover of $\hat{X}$. Denote by $h$ the natural morphism $Z\rightarrow\tilde{X}$. By construction, $Z$ is the pull-back of $EG$ via $c_{\tilde{X},X}\circ f$, so that
$c_{\tilde{X},X}\circ f=c_{Z,\hat{X}}$
is the classifying map of the universal cover $Z$ as a $G$-principal bundle on $\hat{X}$.
Since $[\hat{\omega}^{n-1}]\in V^{2n-2}_{\textrm{hyp}}(\hat{X})$, by Corollary 3.1 there exists a class $[\alpha]\in V^{2n-2}_{\textrm{hyp}}(BG)$ such that $c^{\ast}_{Z,\hat{X}}[\alpha]=[\hat{\omega}^{n-1}]$. Let $[\Omega]=c^{\ast}_{\tilde{X},X}[\alpha]\in V^{2n-2}_{\textrm{hyp}}(X)$. Then we have $[\hat{\omega}^{n-1}]=f^{\ast}[\Omega]$ by construction. Moreover, since $f$ is a small contraction, we also have
$f_{\ast}[\hat{\omega}^{n-1}]=f_{\ast}f^{\ast}[\Omega]=[\Omega].$
Remember that $[\Omega]=f_{\ast}[\hat{\omega}^{n-1}]$ is balanced as is shown before. It is exactly the desired balanced hyperbolic class on $X$.
(2) Assume that $(X,\omega)$ is semi-balanced hyperbolic. It means that $\omega$ is non-negative and strictly positive on a Zariski open set $U$, so $f^{\ast}\omega$ is non-negative and strictly positive on a Zariski open set $f^{-1}(U)\setminus E$. The $\tilde{d}$-boundedness of $f^{\ast}\omega^{n-1}$ is directly by Lemma 2.3. In summary, $f^{\ast}\omega$ is semi-balanced hyperbolic.
(3) Assume that $(\hat{X},\hat{\beta})$ is weakly balanced $k$-hyperbolic. Then $[\hat{\beta}]\in V^{2k}_{\textrm{hyp}}(\hat{X})$. Here we use the fact that the $\tilde{d}$-boundedness of $\hat{\beta}$ implies the $d$-closedness. Similar with the proof of (1), there exists a class $[\alpha]\in V^{2k}_{\textrm{hyp}}(BG)$ such that $c^{\ast}_{Z,\hat{X}}[\alpha]=[\hat{\beta}]$. Let
$[\beta]=c^{\ast}_{\tilde{X},X}[\alpha]\in V^{2k}_{\textrm{hyp}}(X).$
Then $[\hat{\beta}]=c^{\ast}_{Z,\hat{X}}[\alpha]=f^{\ast}c^{\ast}_{\tilde{X},X}[\alpha]=f^{\ast}[\beta]$. Since $f$ is a small contraction, we also have
$f_{\ast}[\hat{\beta}]=f_{\ast}f^{\ast}[\beta]=[\beta].$
By hypothesis $k\mid(n-1)$, say $kt=n-1$. Then
$[\hat{\beta}^{t}]=f^{\ast}[\beta^{t}]\quad\textrm{and}\quad f_{\ast}[\hat{\beta}^{t}]=[\beta^{t}].$
Now we are left to prove the nefness and bigness of $[\Omega:=\beta^{t}]$. Suppose $[\Omega]$ is not nef. By (2.1) there exists a non-zero pseudo-effective $(1,1)$-class $[P]$ on $X$, such that
$[\hat{\beta}^{t}]\cdot f^{\ast}[P]=[\Omega]\cdot[P]<0.$ (3.1)
It contradicts to the fact that $[\hat{\beta}^{t}]$ is nef and $f^{\ast}[P]$ is pseudo-effective. Therefore $[\Omega]$ is nef.
In order to show the bigness, we should apply the duality
$\overline{\mathcal{K}}=\mathcal{N}^{\vee}$ (3.2)
in [BDPP13], Theorem 2.1. Suppose $[\Omega]$ is not big. By (3.2) there exists a non-zero nef $(1,1)$-class $[\gamma]$ on $X$, such that
$[\hat{\beta}^{t}]\cdot f^{\ast}[\gamma]=[\Omega]\cdot[\gamma]<0.$
It contradicts to the fact that $[\hat{\beta}^{t}]$ is big and $f^{\ast}[P]$ is nef. Therefore $[\Omega]$ is big. In summary, we obtain that $[\beta]$ is weakly balanced $k$-hyperbolic. ∎
It is possible to extend the duality (3.2) to the non-Kähler case, hence the Kähler assumption in (3) is not necessary. We will discuss this topic in an upcoming paper. Also one may wonder that if $(X,\beta)$ is weakly balanced $k$-hyperbolic, whether $\hat{X}$ will be. Naturally we could consider $f^{\ast}\beta^{t}$, which surely inherits the nefness and $\tilde{d}$-boundedness. However, as is pointed out in [LX16], Sect.5.7, a big $(n-1,n-1)$-class can pull back to a class on the pseudo-effective boundary. Perhaps this direction is not true.
3.3 The smooth modification
Proof of Theorem 1.2.
(1) Assume that $(\hat{X},\hat{\omega})$ is balanced hyperbolic. The same argument as in the proof of Theorem 1.1 implies that $[f_{\ast}\hat{\omega}^{n-1}]$ is balanced hyperbolic on $X$.
(2) & (3) The proof of Theorem 1.1, (2) and (3) applies verbatim here. ∎
Note that $\hat{X}$ is surely a balanced manifold provided $X$ is by [AB91]. However, the original proof therein doesn’t work for the hyperbolicity. Hence it is still open whether $\hat{X}$ is a balanced hyperbolic manifold provided $X$ is.
4 Improvement
This section aims to prove that the class of balanced hyperbolic manifolds is invariant under the curve contraction, without limitations on the higher homotopy groups any more. In our paper, a curve contraction $f:\hat{X}\rightarrow X$ between compact complex manifolds of dimension $\geqslant 2$ is a holomorphic map such that, for a suitable collection of points $Y$ in $X$, $E:=f^{-1}(Y)$ is a curve and $f|_{\hat{X}\setminus E}:\hat{X}\setminus E\rightarrow X\setminus Y$ is a biholomorphism. The proof utilizes a modification of the topological technique developed in [BKS24] and [BP21].
4.1 Topological preparation II
We first prove a variant of Proposition 2.3 of [BP21].
Lemma 4.1.
Let $f:\hat{X}\rightarrow X$ be a curve contraction of compact complex manifolds of dimension $n$. Then for $31\leqslant i\leqslant 2n-$ 3, $f_{\ast}:\pi_{i}(\hat{X})\rightarrow\pi_{i}(X)$ are isomorphisms.
Proof.
Let $M\subseteq\hat{X}$ and $N\subseteq X$ be the minimal analytic sets so that
$f|_{\hat{X}\setminus M}:\hat{X}\setminus M\rightarrow X\setminus N$
is a biholomorphism. By assumption that $f$ is a curve contraction, it is clear that
$\dim_{\mathbb{C}}M=1\quad\textrm{and}\quad\dim_{\mathbb{C}}N=0.$
Then $N$ is decomposed into finite points as $N=\cup^{l}_{k=1}\{p_{k}\}$, and $M$ also has a decomposition, which is denoted by
$Q_{1},Q_{2}\quad\textrm{with}\quad 0\leqslant\dim_{\mathbb{C}}Q_{1}\leqslant\dim_{\mathbb{C}}Q_{2}=1.$
Notice that $Q_{1}$ and $Q_{1}\cup Q_{2}=M$ are analytic subsets of $\hat{X}$, and $Q_{1}$ and $Q_{2}$ are smooth. Both of these two decompositions naturally satisfy the conditions mentioned at the beginning of [BP21], Proposition 2.3.
Now let $\iota:X\setminus\{p_{1}\}\rightarrow X$ be the natural inclusion. Take a neighbourhood $V$, which is homeomorphic to a $2n$-cell, of $\{p_{1}\}$. Since $(X,X\setminus\{p_{1}\})$ is homotopic to the pair $(X,X\setminus V)$, we can instead consider the long exact sequence as follows:
$\cdots\rightarrow\pi_{k+1}(X,X\setminus V)\rightarrow\pi_{k}(X\setminus V)\overset{\iota_{\ast}}{\rightarrow}\pi_{k}(X)\rightarrow\pi_{k}(X,X\setminus V)\rightarrow\cdots.$
Note $\pi_{k}(X,X\setminus V)=0$ when $k\leqslant 2n-1$. For $21\leqslant k\leqslant 2n-$ 2, $\iota$ induces following isomorphisms
$\iota_{\ast}:\pi_{k}(X\setminus\{p_{1}\})\rightarrow\pi_{k}(X)$
and epimorphism
$\iota_{\ast}:\pi_{2n-1}(X\setminus\{p_{1}\})\rightarrow\pi_{2n-1}(X).$
Inductively, we conclude that the inclusion $j:X\setminus N\rightarrow X$ induces isomorphisms
$j_{\ast}:\pi_{k}(X\setminus N)\rightarrow\pi_{k}(X)\quad\textrm{for }1\leqslant k\leqslant 2n-2.$
Similarly, the inclusion $i:\hat{X}\setminus M\rightarrow\hat{X}$ induces isomorphisms
$i_{\ast}:\pi_{k}(\hat{X}\setminus M)\rightarrow\pi_{k}(\hat{X})\quad\textrm{for }1\leqslant k\leqslant 2n-4,$
and epimorphism $i_{\ast}:\pi_{2n-3}(\hat{X}\setminus M)\rightarrow\pi_{2n-3}(\hat{X})$.
Clearly $(f\circ i)_{\ast}=(j\circ(f|_{\hat{X}\setminus M}))_{\ast}$. Since $f|_{\hat{X}\setminus M}$ is biholomorphic,
$(j\circ(f|_{\hat{X}\setminus M}))_{\ast}:\pi_{k}(\hat{X}\setminus M)\rightarrow\pi_{k}(X)$
are actually isomorphisms for $31\leqslant k\leqslant 2n-$ 3. Consequently, $i_{\ast}:\pi_{2n-3}(\hat{X}\setminus M)\rightarrow\pi_{2n-3}(\hat{X})$ is moreover injective, hence also isomorphic. As a result, for $31\leqslant k\leqslant 2n-$ 3,
$f_{\ast}:\pi_{k}(\hat{X})\rightarrow\pi_{k}(X)$
are isomorphisms. ∎
A simple adjustment of the proof of Lemma 4.1 implies the following result.
Corollary 4.1.
Let $f:\hat{X}\rightarrow X$ be a birational morphism between compact complex manifolds of dimension $n$, such that the exceptional set $E$ satisfies $\dim E=n-k$. Then for $11\leqslant i\leqslant 2k-$ 1, $f_{\ast}:\pi_{i}(\hat{X})\rightarrow\pi_{i}(X)$ are isomophisms.
Now let $X$ be a compact smooth manifold (or more general a finite simplicial complex) of real dimension $2n$, and let $G=\pi_{1}(X)$. We can construct a model analogy to the classifying space $BG$ as follows. Consider the $(2n-3)$-th term $U$ of the Postnikov tower of $X$ by gluing cells of dimension $12n-$ 1 or higher to kill all the $i$-th homotopy groups for $i\geqslant 2n-2$. As we can furthermore attach cells of dimension $3$ or higher to obtain $BG$, $U$ is seen as the subcomplex of $BG$. Moreover, the composition of the inclusions $c_{X,n-1}:X\hookrightarrow U$ and $U\hookrightarrow BG$ is the classifying map of the universal cover of $X$. By construction, the $(2n-2)$-skeleton of $U$ is the same as $X$. So we have
Lemma 4.2.
The homomorphism $c_{X,n-1}^{\ast}:H^{2n-2}(U,\mathbb{R})\rightarrow H^{2n-2}(X,\mathbb{R})$ is injective, and
$c_{X,n-1}^{\ast}(H^{2n-2}(U,\mathbb{R}))=V^{2n-2}_{\textrm{asph}}(X).$
Proof.
As $\pi_{2n-2}(U)=0$, it is clear that $V^{2n-2}_{\textrm{asph}}(U)=H^{2n-2}(U,\mathbb{R})$. Hence by definition
$c_{X,n-1}^{\ast}(H^{2n-2}(U,\mathbb{R}))\subseteq V^{2n-2}_{\textrm{asph}}(X).$
Now consider the the long exact sequence of cohomology associated to $(X,\,U)$:
$\cdots\rightarrow H^{2n-2}(U,X;\mathbb{R})\rightarrow H^{2n-2}(U,\mathbb{R})\xrightarrow{c_{X,n-1}^{\ast}}H^{2n-2}(X,\mathbb{R})\xrightarrow{\partial}H^{2n-1}(U,X;\mathbb{R})\rightarrow\cdots.$
By the universal coefficient theorem for relative cohomology and the fact that $X$ and $U$ share the same $(2n-2)$-skeleton, the relative cohomology group $H^{2n-2}(U,X;\mathbb{R})$ vanishes and $c_{X,n-1}^{\ast}$ is injective.
It is left to show $V^{2n-2}_{\textrm{asph}}(X)\subseteq\text{Ker}\,\partial$. Let $[\omega]\in V^{2n-2}_{\textrm{asph}}(X)$. Then $\partial[\omega]\in H_{2n-1}(U,X;\mathbb{R})^{\ast}$ acts on a relative $(2n-1)$-cycle $\gamma+C_{2n-1}(X)$ by
$\partial[\omega](\gamma+C_{2n-1}(X))=\omega(\partial_{2n-1}\gamma),$
where $\partial_{2n-1}$ is boundary operator. Notice that $\partial_{2n-1}\gamma\in Z_{2n-2}(X)$ which, if non-trivial, comes from a $(2n-1)$-cell attached to $X$. Thus $\partial_{2n-1}\gamma$ is combination of the images of $S^{2n-2}$. As $\omega$ is a aspherical class, it follows that $\omega(\partial_{2n-1}\gamma)=0$. Namely $[\omega]\in\text{Ker}\,\partial$. ∎
If we instead construct the $(2k-1)$-th term $U_{k}$ of the Postnikov tower of $X$ by gluing cells of dimension $12k+$ 1 or higher to kill all the $i$-th homotopy groups for $i\geqslant 2k$, and let $c_{X,k}:X\hookrightarrow U_{k}$ be the inclusion, then the same argument implies that
Corollary 4.2.
The homomorphism $c_{X,k}^{\ast}:H^{2k}(U_{k},\mathbb{R})\rightarrow H^{2k}(X,\mathbb{R})$ is injective, and
$c_{X,k}^{\ast}(H^{2k}(U_{k},\mathbb{R}))=V^{2k}_{\textrm{asph}}(X).$
Next we show that $U$ possesses the same universal property for the hyperbolic $(2n-2)$-classes as $BG$.
Proposition 4.1.
Let $X$ and $Y$ be two finite simplicial complexes. Let $f:Y\rightarrow U$ be a continuous map, and $[\omega]\in H^{2n-2}(U,\mathbb{R})$. If $c_{X,n-1}^{\ast}[\omega]\in V^{2n-2}_{\textrm{hyp}}(X)$, then $f^{\ast}[\omega]\in V^{2n-2}_{\textrm{hyp}}(Y)$.
Proof.
As $X\subseteq U$ and $Y$ are finite complexes, there is a finite subcomplex $X^{\prime}$ of $U$, so that both $X\subseteq X^{\prime}$ and $f(Y)\subseteq X^{\prime}$. Remember the hyperbolicity is preserved under the pull-back of a simplicial map and every continuous map is approximated by simplicial ones. It is sufficient to prove that $[\omega|_{X^{\prime}}]$ is a hyperbolic class on $X^{\prime}$.
Notice that $X^{\prime}$ is obtained by gluing finite cells of dimension $(2n-1)$ or higher to $X$. Assume first the number of glued cells equals one, and denote the gluing map by
$h:S^{k-1}\rightarrow X\quad\textrm{with }k\geqslant 2n-1.$
Then $X^{\prime}=X\cup_{h}D^{k}$. Let $\pi:\tilde{X}\rightarrow X$ be the universal cover of $X$, then the universal cover of $X^{\prime}$ is just $\tilde{X}^{\prime}=\tilde{X}\cup_{h\times\pi}(D^{k}\times\pi_{1}(X))$. Let $\pi:\tilde{X}^{\prime}\rightarrow X^{\prime}$ be the natural projection by abuse the notation. Choose a representative $\omega$ of $[\omega]$ in $\Omega^{2n-2}(U)$. As $c_{X,n-1}^{\ast}\omega$ is hyperbolic on $X$, by definition, there is a bounded $(2n-3)$-form $\alpha$ on $\tilde{X}$, so that $d\alpha=\pi^{\ast}(\omega|_{X})$. Now consider $\omega|_{D^{k}}$. As $H^{2n-2}_{\textrm{dR}}(D^{k})=0$ for $k\geqslant 2n-1$, there is an $\alpha^{\prime}\in\Omega^{2n-3}(D^{k})$, so that $d\alpha^{\prime}=\omega|_{D^{k}}$. In particular, $\alpha^{\prime}$ is also bounded since $D^{k}$ is compact. Let $h_{0}$ (resp. $\mathbb{D}$) be an arbitrary lift of $h$ (resp. $D^{k}$). Denote by $S^{k-1}_{0}$ the boundary of one sheet $D^{k}_{0}$ of $\mathbb{D}$. Then by the choice of $h$,
$d(h^{\ast}_{0}\alpha|_{S^{k-1}_{0}})=(h^{\ast}_{0}d\alpha)|_{S^{k-1}_{0}}=h^{\ast}_{0}(\pi^{\ast}\omega)|_{S^{k-1}_{0}},$
where the naturality of $d$ is ensured by De Rham-Thom theorem for finite simplicial complexes. Thus
$d((h^{\ast}_{0}\alpha-h^{\ast}_{0}\pi^{\ast}\alpha^{\prime})|_{S^{k-1}_{0}})=h^{\ast}_{0}(\pi^{\ast}\omega)|_{S^{k-1}_{0}}-h^{\ast}_{0}(\pi^{\ast}\omega)|_{S^{k-1}_{0}}=0.$
It means that $(h^{\ast}_{0}\alpha-h^{\ast}_{0}\pi^{\ast}\alpha^{\prime})|_{S^{k-1}_{0}}\in\text{Ker}\,d$. As $k\geqslant 2n-1$, $H^{2n-3}_{\textrm{dR}}(S^{k-1}_{0})=0$. Thus there exists a $\xi_{0}\in\Omega^{2n-4}(S^{k-1}_{0})$, such that $d\xi_{0}=(h^{\ast}_{0}\alpha-h^{\ast}_{0}\pi^{\ast}\alpha^{\prime})|_{S^{k-1}_{0}}$. The boundedness of $d\xi_{0}$ is ensured by the boundedness of $\alpha$ and $\alpha^{\prime}$. Pick a smooth extension $\xi^{\prime}_{0}$ of $\xi_{0}$ on $D^{k}_{0}$, so that $d\xi^{\prime}_{0}$ is still bounded. Now we can patch $\alpha$ and $\pi^{\ast}\alpha^{\prime}+d\xi^{\prime}_{0}$ together to obtain an $\hat{\alpha}_{0}$ on $\tilde{X}\cup_{h_{0}}D^{k}_{0}$ such that $\pi^{\ast}\omega=d\hat{\alpha}_{0}$ restricted on this space. Following this streamline, we obtain an $\hat{\alpha}$ on the whole $\tilde{X}^{\prime}$ such that $\pi^{\ast}\omega=d\hat{\alpha}$. Since $\hat{\alpha}$ is bounded, $[\omega|_{X^{\prime}}]$ is hyperbolic.
When the number of glued cells is larger, we inductively obtain the desired result. ∎
It naturally extends to the following form.
Corollary 4.3.
Let $X$ and $Y$ be two finite simplicial complexes. Let $f:Y\rightarrow U_{k}$ be a continuous map, and $[\omega]\in H^{2k}(U_{k},\mathbb{R})$. If $c_{X,k}^{\ast}[\omega]\in V^{2k}_{\textrm{hyp}}(X)$, then $f^{\ast}[\omega]\in V^{2k}_{\textrm{hyp}}(Y)$.
Define the subspace $V^{k}_{\textrm{hyp}}(U)$ of the real singular cohomology group $H^{k}(U,\mathbb{R})$ to be the set of $k$-cohomology classes whose pull-back to any finite simplicial complex is hyperbolic. We have
Proposition 4.2.
$c_{X,n-1}^{\ast}(V^{2n-2}_{\textrm{hyp}}(U))=V^{2n-2}_{\textrm{hyp}}(X)$.
Proof.
By definition, $c_{X,n-1}^{\ast}(V^{2n-2}_{\textrm{hyp}}(U))\subseteq V^{2n-2}_{\textrm{hyp}}(X)$. On the other hand, if $[\omega]\in V^{2n-2}_{\textrm{hyp}}(X)$, then naturally $[\omega]\in V^{2n-2}_{\textrm{asph}}(X)$. By Lemma 4.2, there exists an $[\alpha]\in H^{2n-2}(U,\mathbb{R})$, so that $[\omega]=c_{X,n-1}^{\ast}[\alpha]$. On the other hand, $[\omega]$ is hyperbolic, by Proposition 4.1 the continuous pull-back of $[\alpha]$ to any finite simplicial complex is hyperbolic. Therefore $[\alpha]\in V^{2n-2}_{\textrm{hyp}}(U)$. ∎
Define the subspace $V^{i}_{\textrm{hyp}}(U_{k})$ of the real singular cohomology group $H^{i}(U_{k},\mathbb{R})$ to be the set of $i$-cohomology classes whose pull-back to any finite simplicial complex is hyperbolic. Replacing Lemma 4.2 and Proposition 4.1 by Corollaries 4.2 and 4.3, we obtain
Corollary 4.4.
$c_{X,k}^{\ast}(V^{2k}_{\textrm{hyp}}(U_{k}))=V^{2k}_{\textrm{hyp}}(X)$.
4.2 The curve contraction
We are in a good position to prove Theorems 1.3 and 1.4. In particular, the vanishing requirement for the higher homotopy groups is not necessary any more.
Proof of Theorem 1.4.
(1) Assume that $(\hat{X},\hat{\Omega})$ is balanced hyperbolic. Let $G=\pi_{1}(X)$, and $\tilde{X}$ is the universal covering of $X$. Keep the notations before, we should consider the following commutative diagram:
${U}$${\hat{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{c_{\hat{X},1}\simeq c_{X,1}\circ f}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{c_{X,1}}$
By Lemma 4.1, $f_{\ast}:\pi_{1}(\hat{X})\rightarrow\pi_{1}(X)$ and $(c_{X,1})_{\ast}:\pi_{1}(X)\rightarrow\pi_{1}(U)$ are isomorphisms, while for $i\geq 2$, $\pi_{i}(U)=0$. Then by the uniqueness (up to the homotopy) of the Postnikov tower of $\hat{X}$ (see [Hat02]), $c_{\hat{X},1}:\hat{X}\rightarrow U$ is the first term of the Postnikov tower of $\hat{X}$, and we have the homotopy equivalence $c_{X,1}\circ f\simeq c_{\hat{X},1}$. Moreover, as $U$ can be regarded as a subcomplex of $BG$, the universal cover $Z:=f^{\ast}\tilde{X}$ of $\hat{X}$, is actually the pull-back of $EG|_{U}$ by $c_{X,1}\circ f$.
As $[\hat{\Omega}]\in V^{2}_{\textrm{hyp}}(\hat{X})$, by Proposition 4.2 there exists a unique $[\alpha]\in V^{2}_{\textrm{hyp}}(U)$, so that $c_{\hat{X},1}^{\ast}[\alpha]=[\hat{\Omega}]$. Denote $[\beta]=c_{X,1}^{\ast}[\alpha]\in V^{2n-2}_{\textrm{hyp}}(X)$, thus $[\hat{\Omega}]=f^{\ast}[\beta]$. Moreover, as $f$ is a curve contraction,
$f_{\ast}[\hat{\Omega}]=f_{\ast}f^{\ast}[\beta]=[\beta].$
Remember that $[\beta]=f_{\ast}[\hat{\Omega}]$ is balanced as is shown before. It is exactly the desired balanced hyperbolic class on $X$.
(2) Assume that $(X,\Omega)$ is balanced hyperbolic. As is shown in Sect.3.2, the real $d$-closed $(n-1,n-1)$-class $f^{\ast}[\Omega]$ is balanced. (When $\dim X>2$, a curve contraction must be a small contraction!) Hence there exists a balanced metric $\hat{\Omega}$ such that $[\hat{\Omega}]=f^{\ast}[\Omega]$. On the other hand, since $\Omega$ is $\tilde{d}$-bounded, so is $f^{\ast}\Omega$ by Lemma 2.3. It implies that $\hat{\Omega}$ is $\tilde{d}$-bounded again by Lemma 2.3.
Conversely, assume that $(\hat{X},\hat{\Omega})$ is balanced hyperbolic. Let $G=\pi_{1}(X)$, and $\tilde{X}$ is the universal covering of $X$. Keep the notations before, we should consider the following commutative diagram:
${U}$${\hat{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{c_{\hat{X},n-1}\simeq c_{X,n-1}\circ f}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{c_{X,n-1}}$
By Lemma 4.1, for $31\leqslant i\leqslant 2n-$ 3, $f_{\ast}:\pi_{i}(\hat{X})\rightarrow\pi_{i}(X)$ and $(c_{X,n-1})_{\ast}:\pi_{i}(X)\rightarrow\pi_{i}(U)$ are isomorphisms, while for $i\geq 2n-2$, $\pi_{i}(U)=0$. Then by the uniqueness (up to the homotopy) of the Postnikov tower of $\hat{X}$ (see [Hat02]), $c_{\hat{X},n-1}:\hat{X}\rightarrow U$ is the $(2n-3)$-th term of the Postnikov tower of $\hat{X}$, and we have the homotopy equivalence $c_{X,n-1}\circ f\simeq c_{\hat{X},n-1}$. Moreover, as $U$ can be regarded as a subcomplex of $BG$, the universal cover $Z:=f^{\ast}\tilde{X}$ of $\hat{X}$, is actually the pull-back of $EG|_{U}$ by $c_{X,n-1}\circ f$.
As $[\hat{\Omega}]\in V^{2n-2}_{\textrm{hyp}}(\hat{X})$, by Proposition 4.2 there exists a unique $[\alpha]\in V^{2n-2}_{\textrm{hyp}}(U)$, so that $c_{\hat{X},n-1}^{\ast}[\alpha]=[\hat{\Omega}]$. Denote $[\beta]=c_{X,n-1}^{\ast}[\alpha]\in V^{2n-2}_{\textrm{hyp}}(X)$, thus $[\hat{\Omega}]=f^{\ast}[\beta]$. Moreover, as $f$ is a curve contraction,
$f_{\ast}[\hat{\Omega}]=f_{\ast}f^{\ast}[\beta]=[\beta].$
Remember that $[\beta]=f_{\ast}[\hat{\Omega}]$ is balanced as is shown before. It is exactly the desired balanced hyperbolic class on $X$.
(3) Assume $(\hat{X},\hat{\Omega})$ is weakly balanced $(n-1)$-hyperbolic. Apply the same argument in (1) and keep the notations there, we obtain a unique $[\alpha]\in V^{2n-2}_{\textrm{hyp}}(U)$, so that $c_{\hat{X},n-1}^{\ast}[\alpha]=[\hat{\Omega}]$. Denote $[\beta]=c_{X,n-1}^{\ast}[\alpha]\in V^{2n-2}_{\textrm{hyp}}(X)$, thus $[\hat{\Omega}]=f^{\ast}[\beta]$. Moreover, as $f$ is a curve contraction,
$f_{\ast}[\hat{\Omega}]=f_{\ast}f^{\ast}[\beta]=[\beta].$
Now a similar application of the cone duality as in the proof of Theorem 1.1, (3) implies that $[\beta]$ is nef and big. In summary, $(X,\beta)$ is weakly balanced $(n-1)$-hyperbolic. ∎
Since on a complex surface, a balanced metric is equivalent to a Kähler metric, (1) is reformulated as, if $\hat{X}$ is Kähler hyperbolic, so will be $X$.
Proof of Theorem 1.4.
Assume $(\hat{X},\hat{\beta})$ is weakly balanced $k$-hyperbolic. If we replace Lemma 4.1 and Proposition 4.2 by Corollaries 4.1 and 4.4, the proof of Theorem 1.4 applies verbatim here. In conclusion there exists a unique $[\alpha]\in V^{2k}_{\textrm{hyp}}(U_{k})$, so that $c_{\hat{X},k}^{\ast}[\alpha]=[\hat{\beta}]$. Denote
$[\beta]=c_{X,k}^{\ast}[\alpha]\in V^{2k}_{\textrm{hyp}}(X),$
thus $[\hat{\beta}]=f^{\ast}[\beta]$. Moreover, as $f$ is a curve contraction,
$f_{\ast}[\hat{\beta}]=f_{\ast}f^{\ast}[\beta]=[\beta].$
By hypothesis $k\mid(n-1)$, say $kt=n-1$. Then we have
$[\hat{\beta}^{t}]=f^{\ast}[\beta^{t}]\quad\textrm{and}\quad f_{\ast}[\hat{\beta}^{t}]=[\beta^{t}].$
A similar application of the cone duality as in the proof of Theorem 1.1, (3) implies that $[\beta^{t}]$ is nef and big. In summary, $(X,\beta)$ is weakly balanced $k$-hyperbolic. ∎